# Lasso Multinomial Performance Indicators for in-play Basketball Data

### v1

@@ 2.3.3 Filtering low-time players (LTPs)
Although results improved in both approaches when using the filtered dataset, the ridge-based RAPMs did not show sufficient improvement, as they tend to prioritise unexpectedly performed players at the top of their rankings. For instance, only $28\%$ of the top 100 offensive and defensive ridge RAPMs consisted of starter players. This marks a substantial improvement compared to the corresponding percentage of $8\%$ when all players were included in the analysis. However, this improvement does not appear to be adequate.
@@ 2.3.4 Logistic regression with shrinkage methods
In this section, we proceed by considering as a response variable $Y_{i}$ which records whether the team in possession scored or not. Since this simplified response is now binary, we will consider a standard, Bernoulli-based, logistic regression model given by
@@
For the estimation of logistic regression RAPMs, we have similarly applied regularization methods as in the normal-based RAPMs. For both ridge and lasso methods, the RAPM-based ranking where obtained using the shrinkage parameter value $\lambda_{min}$ obtained from a 10-fold cross-validation. The lasso implementation resulted in 200 player RAPMs (either offensive or defensive) which were shrunk to zero. This represented approximately the $20\%$ of the total estimated RAPMs (and players, respectively).
From Binomial analysis, the correlation between RAPM ratings derived from the ridge Binomial ($\text{RAPM}^{\text{B}}$) and ridge Normal ($\text{RAPM}^{\text{N}}$) methodologies reveals an interesting result. Specifically, from our analysis, we found a significant linear relationship between these two RAPM ratings, as illustrated in Figure 3, expressed by the following equation:
$\text{RAPM}^{\text{B}}_{k}=0.02\times\text{RAPM}^{\text{N}}_{k}+\varepsilon_{k},\quad\varepsilon_{k}\sim\mathcal{N}(0,0.008^{2})$ (3)
for all $k=1,\ldots,K$ players (offender or defender), with $R^{2}\approx 0.89$. This equation allows us to compute the Binomial RAPMs from those derived under the Normal distribution. This capability enhances the interpretability of Normal RAPMs by providing additional insights. Moreover, it provides a reasonable theoretical justification for using the normal RAPMs as approximations of the binomial RAPMs, which are derived from a model which is properly defined with a correctly specified distribution.
Figure 3: Coefficients of ridge Binomial vs ridge Normal model.
@@ 2.3.5 Multinomial logistic regression with shrinkage methods
In this section, we proceed with the implementation of a multinomial logistic regression for modelling the scoring outcome of each possession. As we have already discussed, this approach is more appropriate for our problem since it can model separately each of the types of scores. In this way it accounts for different scoring contributions of each player.
@@
From the above-fitted models, we obtain the multinomial probabilities using Eq. 5. Note that the estimates based on the separate logistic regression approach are less efficient with larger standard errors than the direct multinomial logistic approach. Nevertheless, the loss is minor when the baseline category is dominant on the data like in our case [1, Chap. 8]. We prefer the separate logistic regression approach here for two reasons. First, the separate regression approach offers advantages in terms of computational efficiency. Hence, it can be implemented for datasets with large sample sizes, as in the present study. Secondly, this approach allows us to implement different variable selection and shrinkage algorithms on the effects of each scoring category. By this way, we can identify the contribution of each player in the three different types of scores.
@@
Figure 4: Relative frequencies of actual points per possession vs. the simulated by the lasso Multinomial model.
3 Results
This section presents the key findings obtained by applying the methodological procedures outlined earlier in this work. Our focus lies on comparing the RAPM approaches presented in previous sections. We demonstrate the superiority of the multinomial lasso-RAPMs (EPTS and WEPTS) compared to the ones obtained with ridge and lasso normal models. Prior to this comparison, we establish the external validation criteria that serve as the basis for evaluating the performance of the different approaches.
@@ 3.1 External Validation Criteria
    Top-50 Box Score Statistics Criterion: This criterion identifies the top-50 players based on a variety of box score statistics (points, assists, and offensive rebounds for offence; defensive rebounds, steals, and blocks for defense). We then measure the percentage of these players who are included in the top-50 RAPM rankings.
3.2 Ridge vs. Lasso RAPM Normal Metrics
@@
An examination of Tables 1 and 2 reveals the superiority of lasso over ridge. Specifically, focusing on the playing time of top-rated players (Criterion 2), as shown in Table 1, our analysis highlights a clear advantage of lasso RAPMs over ridge. We observe that 14% of the top 100 players in the ridge RAPM rankings are among those with the lowest recorded playing time. This percentage drops to 3% for lasso Normal and 6% for the OLS ratings after removing the zero-contributed players indicated by lasso. The latter will be referred to as after-lasso RAPMs. This suggests a more accurate performance of the lasso models since it is generally unexpected for players with minimal in-season playtime to appear in the list with the highest contributions. This is a common behaviour of the plus/minus ratings found in ridge RAPMs, but restricted in the lasso approach.
@@
3.3 Lasso Normal vs. Multinomial
After demonstrating the superiority of the lasso technique over ridge regression in Section 3.2, we proceed here to examine the normal and multinomial RAPMs inside the lasso framework. As discussed in Section 2.3.5, the proposed multinomial model accounts for the discrete nature of points scored per possession. This choice, in combination with lasso and its shrinkage properties, offers a clear advantage over the commonly used ridge regression RAPMs. Therefore, in this section, we will examine and compare the external validation criteria (see Section 3.1) for both the lasso multinomial and Normal models.
From Tables  3 and 4 we observe small differences in the performance, with no clear winner between the two distributions for the different evaluation criteria. The lasso multinomial performs better in two out of four evaluation criteria presented in Table 3. When analyzing players from “All NBA teams” (Criterion 1), lasso multinomial outperforms the corresponding normal model by identifying a larger proportion of such players ($47\%$ vs. $40\%$). Furthermore, in Criterion 3b, the multinomial lasso performs marginally better than the standard lasso since fewer starters are included in the bottom 100 list ($20\%$ vs $23\%$). On the other hand, Normal-based RAPMs outperform the multinomial ones in Criterion 2, since a lower proportion of low-time players ($3\%$ vs. $9\%$) is included in its top 100 player rankings. Finally, regarding Criterion 3a, a higher proportion of starters are included in the top-100 RAPM list ($55\%$ vs. $47\%$). Regarding the box-score statistics (Criterion 4) of top-rated players, lasso multinomial is marginally better than normal lasso in all statistics except the number of points (PTS); see Table 4.
Table 3: Comparison of validation Criteria 1–3 between ratings for the lasso Normal and Multinomial models.
@@
Table 4: Comparison of External Validation Criterion 4 for Lasso Ratings of Normal and Multinomial.
@@
From the previous comparisons, we conclude that the lasso multinomial performs somewhat better than the standard approach. Furthermore, the multinomial model is more realistic than the standard normal approach since the response is properly considered as a discrete random variable, whereas the latter approach (incorrectly) assumes that the number of points per possession is a continuous random variable. Finally, the choice of the multinomial model is further supported by the fact that the RAPMs in conventional lasso models are shrunk to zero for around $70\%$ of the players. In contrast, the corresponding proportion for the lasso multinomial model is reduced to only $30\%$. Since the goal is to develop an evaluation metric, reducing the RAPM for the majority of players to zero implies that, for a large number of players, no clear evaluation and discrimination from the “average” level will be provided.
@@ 4.1 The metric
In this section, we proceed by introducing an improved weighted version of EPTS which takes into consideration the participation of each player throughout the season. The goal of this evaluation metric is to quantify the expected points per possession for the team of the player under study. We will evaluate two different lineup configurations: one lineup including the player of interest and a second one without the player of interest. All other players on the court are assumed to belong to the reference category. The corresponding expected points of these two line-ups will be weighted according to weighted according to the proportion of possessions in which the player of interest participated.
@@ 4.2 Weighted EPTS (wEPTS) through validation criteria
The RAPM and EPTS ratings presented on the previous sections are based on biased “shrunk” estimates of model coefficients which quantify the contribution of each player. This section explores the impact of bias on RAPM and EPTS ratings. We compare two approaches: (a) using directly biased coefficients derived using lasso and (b) using the unbiased (MLE) coefficients, after removing players whose coefficients were flagged as zero by lasso. For the latter case we will use the conventional name “after-lasso” ratings.
@@
5 Highlights and Discussion
Following the literature, we initially implemented ridge regression on the full possession dataset for the 2021-2022 NBA season. However, this approach was susceptible to the influence of low-time players (LTPs), whose limited playing time inflated their estimated contributions. To address this issue, we have investigated two different approaches: First, we have implemented lasso instead of ridge and, second, we have removed low-time players from RAPM estimation.
For the first approach, lasso regression promotes sparsity by setting specific coefficients equal to zero. This will be effective for less influential variables (players). This characteristic of lasso resulted in RAPM estimates with a clearer distinction between well-performed, average and low-performed players. Concerning LTPs, the problem was mitigated (but not diminished) by shrinking the coefficients of such players of them to zero (i.e. to the group of “average” players). On the other hand, the exclusion of LTPs (defined as less than 200 minutes for the season) also improved the performance of the RAPM ratings. For this reason, we have proceeded by combining the two strategies (i.e. implementing lasso without LTPs).

### v2

@@ 3.3.3 Filtering low-time players (LTPs)
Although results improved in both approaches when using the filtered dataset, the ridge-based RAPMs did not show sufficient improvement, as they tend to prioritize unexpectedly performed players at the top of their rankings. For instance, only $28\%$ of the top 100 offensive and defensive ridge RAPMs consisted of starter players. This marks a substantial improvement compared to the corresponding percentage of $8\%$ when all players were included in the analysis. However, this improvement does not appear to be adequate.
@@ 3.3.4 Logistic regression with shrinkage methods
In this section, we proceed by considering the response variable $Y_{i}$, which records whether the team in possession scored or not. Since this simplified response is now binary, we will consider a standard, Bernoulli-based, logistic regression model given by
@@
For the estimation of logistic regression RAPMs, we have similarly applied regularization methods as in the normal-based RAPMs. For both ridge and lasso methods, the RAPM-based ranking where obtained using the shrinkage parameter value $\lambda_{min}$ obtained from a 10-fold cross-validation. The lasso implementation resulted in 617 player RAPMs (either offensive or defensive) which were shrunk to zero. This represented approximately the $63.8\%$ of the total estimated RAPMs (and players, respectively).
[…]
@@ 3.3.5 Multinomial logistic regression with shrinkage methods
In this section, we proceed with the implementation of a multinomial logistic regression for modelling the scoring outcome of each possession. As we have already discussed, this approach is more appropriate for our problem since it can model separately each of the types of scores. In this way it accounts for the different scoring contributions of each player.
@@
From the above-fitted models, we obtain the multinomial probabilities using (4). Note that the estimates based on the separate logistic regression approach are less efficient with larger standard errors than the direct multinomial logistic approach. Nevertheless, the loss is minor when the baseline category is dominant on the data like in our case [1, Chap. 8]. We prefer the separate logistic regression approach here for two reasons. First, the separate regression approach offers advantages in terms of computational efficiency. Hence, it can be implemented for datasets with large sample sizes, as in the present study. Secondly, this approach allows us to implement different variable selection and shrinkage algorithms on the effects of each scoring category. In this way, we can identify the contribution of each player in the three different types of scores.
@@
For instance, when comparing two players, A and B, we consider the following statements:
  • In the scenario where $EPTS^{o}_{A}=0.9$ and $EPTS^{o}_{B}=1$, Player B contributes one (1) point per possession to his team’s offensive performance when he is on the court, which is higher than Player A’s contribution of 0.9 points per possession, considering that all other players on the court are the same. This indicates that Player B has a greater offensive impact, as his presence in the lineup results in a more favorable scoring outcome for his team compared to the corresponding scoring outcome for Player A.
  • Regarding the defensive EPTS measures, let us consider the case where $EPTS^{d}_{A}=0.9$ and $EPTS^{d}_{B}=1$. Then, Player B’s defensive contribution is measured as one point conceded per possession when he is not playing. This is compared to 0.9 points conceded per possession when Player A is not participating in the specific possession, assuming that all other players on the court remain the same. This indicates that Player B’s absence has a greater impact on the team’s defense, as the team concedes a higher number of points per possession when he is not on the court. Consequently, this suggests that Player B is more valuable to his team’s defense than Player A.
By defining EPTS as described above, player comparisons in both offensive and defensive aspects of the game become straightforward and easy to interpret, by using a single metric for each case rather than three separate ones for the probabilities of scoring or conceding one, two, three, or more points. Moreover, the interpretation of offensive and defensive ratings is consistent: a higher EPTS score always signifies a greater contribution in both scenarios. Therefore, the higher the EPTS value, the greater the player’s (offensive or defensive) impact on his team.
@@
Figure 2: Relative frequencies of actual points per possession vs. the fitted ones estimated by the lasso multinomial model using simulation.
4 Results and Comparisons
This section presents the key findings obtained by applying the methodological procedures outlined earlier in this work. Our focus lies in comparing the RAPM approaches presented in previous sections. We demonstrate the superiority of the multinomial lasso-RAPMs (EPTS and WEPTS) compared to the ones obtained with ridge and lasso normal models. Prior to this comparison, we establish the external validation criteria that serve as the basis for evaluating the performance of the different approaches.
@@ 4.1 External Validation Criteria
    Top-50 Box Score Statistics Criterion: This criterion identifies the top-50 players based on a variety of box score statistics (points, assists, and offensive rebounds for offence; defensive rebounds, steals, and blocks for defence). We then measure the percentage of these players who are included in the top-50 RAPM rankings.
4.2 Lasso vs. Ridge RAPMs in Normal Models
@@
An examination of Tables 1 and 2 reveals the superiority of lasso over ridge. Specifically, focusing on the playing time of top-rated players (Criterion 2), as shown in Table 1, our analysis highlights a clear advantage of lasso RAPMs over ridge ones. We observe that 14% of the top 100 players in the ridge RAPM rankings are among those with the lowest recorded playing time. This percentage drops to 3% for lasso normal and 6% for the OLS ratings after removing the zero-contributed players indicated by lasso. The latter will be referred to as after-lasso RAPMs. This suggests a more accurate performance of the lasso models since it is generally unexpected for players with minimal in-season playtime to appear in the list with the highest contributions. This is a common behaviour of the plus-minus ratings found in ridge RAPMs, but restricted in the lasso approach.
@@
4.3 Lasso Multinomial vs. Benchmarks
After demonstrating the superiority of the lasso technique over ridge regression in Section 4.2, we will now examine the proposed multinomial model, compared with the RAPMs obtained by two “benchmark” models: (a) the lasso-normal and (b) the lasso-logistic.
As discussed in Section 3.3.5, the proposed multinomial model accounts for the discrete nature of points scored per possession. This choice, in combination with lasso and its shrinkage properties, offers a clear advantage over the commonly used ridge regression RAPMs. Therefore, in this section, we will examine and compare the external validation criteria (see Section 4.1) for the the three models under consideration.
From Tables 3 and 4, we observe small differences in the performance, with no clear winner between the opposite distributions for the different evaluation criteria. The lasso multinomial performs better in two out of four evaluation criteria presented in Table 3. When analyzing players from “All NBA teams” (Criterion 1), lasso multinomial outperforms both benchmark models by identifying a larger proportion of such players; $33\%<40\%<47\%$ for binomial, normal and multinomial, respectively. Furthermore, in Criterion 3b, the multinomial lasso performs marginally better than the rest cases, since fewer starters are included in the bottom 100 list ($20\%<23\%<44\%$). On the other hand, normal-based RAPMs outperform in Criterion 2, since a lower proportion of low-time players ($3\%$) is included in its top 100 player rankings. Finally, regarding Criterion 3a, a higher proportion of starters are included in the top-100 normal RAPM list ($55\%$). Regarding the box-score statistics (Criterion 4) of top-rated players, lasso multinomial is marginally better than normal lasso in all statistics except the number of points (PTS). In terms of differences in box score statistics between the multinomial and logistic regression models, the proposed model outperforms the binomial one in two criteria: offensive rebounds and steals, while the binomial model is better only in defensive rebounds; see Table 4.
Table 3: Comparison of validation Criteria 1–3 between ratings for the lasso Normal, Binomial and Multinomial models.
@@
Table 4: Comparison of External Validation Criterion 4 for lasso ratings of Normal, Binomial and Multinomial.
@@
Based on the previous comparisons, we conclude that the lasso multinomial model performs somewhat better than the standard normal and binomial models. Furthermore, the multinomial model is more realistic than the standard normal approach since the response is properly considered as a discrete random variable, whereas the normal approach (incorrectly) assumes that the number of points per possession is a continuous random variable. In the case of the binomial model, the different types of points scored are not taken into consideration, which ignores important information about the basketball game. Finally, the choice of the multinomial model is further supported by the fact that the RAPMs in conventional lasso models are shrunk to zero for around $70.0\%$, and $63.8\%$ of the players in the normal, and binomial case, respectively. In contrast, the corresponding proportion for the lasso multinomial model is reduced to only $30\%$. Since the goal is to develop an evaluation metric, reducing the RAPM for the majority of players to zero implies that, for a large number of players, no clear evaluation and discrimination from the “average” level will be provided.
4.4 Joint vs. Separate estimation
As discussed by [12], joint estimation in multinomial logistic regression is generally more efficient than the separate logistic regression approach because it takes into consideration the entire dataset across all outcome categories, leading to more precise and stable parameter estimates. This approach, however, leads to a less flexible model specification, as it assumes the same sets of covariates across all categories.
On the other hand, separate estimation offers greater flexibility by allowing for different relationships between predictors and each outcome category. This flexibility, while advantageous in certain scenarios, can result in higher variance in the estimates due to the reduced amount of data used in each independent model. Nonetheless, as noted by [1], this loss of efficiency is lower when the baseline category is dominant within the data, as is the case in our analysis. In such scenarios, the impact of larger standard errors is negligible, making the separate logistic regression approach a realistic alternative to the multinomial model, particularly when minimizing the computational burden is set as a priority.
For the joint multinomial implementation, a slightly higher number of players effect is shrunk to zero than the separate estimation (305 vs. 292; 31.4% vs. 30%). Moreover, the relationship between the Expected Points per Possession derived from the joint and separate approach is strong (Pearson’s $r\approx 95\%$) while no significant differences are observed in the ECDF and distribution plots; see Figures 5 and 6 respectively.
Figure 5: Empirical cumulative distribution function (ECDF) for the lasso EPTS and wEPTS obtained by the joint and separate estimation of the multinomial model.
Figure 6: Distribution of the lasso EPTS and wEPTS obtained by the joint and separate estimation of the multinomial model.
Based on the results, it seems that the simplified EPTS ratings derived by separate multinomial regression approach offers an accurate and less demanding alternative to EPTS derived by the joint multinomial model approach.
4.5 Poisson implementation
Considering the discrete numerical nature of the response data, an additional candidate model for the specific analysis, is the Poisson regression, as also suggested by a referee. In our context, the Poisson regression model is given by:
$\text{Pts}_{i}\sim\text{Poisson}(\xi_{i})\quad$ (11)
where $\text{Pts}_{i}$ is the number of points scored in possession $i$, $\xi_{i}$ is the expected number of points scored in possession $i$, and $\log(\xi_{i})$ has the same form as $\mu_{i}$ in (1). Regarding the regularization implementation of the Poisson model, the lasso is used as in the previous implementations. Using lasso for Poisson regression models is common in football; see for example in [10] who implemented $L_{1}$-penalized Poisson regression to predict outcomes in international soccer tournaments.
An important limitation of the Poisson model is that the response variable in our case study (i.e., the number of points per possession) is not defined over the entire set of natural numbers, as would be expected for a Poisson-distributed variable. Additionally, a key distinction between the two models lies in their interpretations: the Poisson model focuses on modeling the overall average number of points scored per possession, which indirectly determines the scoring probabilities, while the multinomial model directly estimates each player’s effect on individual scoring probabilities. As a consequence, in a Poisson model, an increased average contribution leads to higher probabilities of scoring two or three points. However, this is not typical of how points are distributed in basketball. Figure 7 illustrates a comparison between the observed and Poisson-fitted probabilities in our dataset. The probabilities of the Poisson model shows a decreasing pattern, whereas the observed probabilities follow a non-decreasing pattern, with the highest probability for no points, followed by two points, three points, and lastly, one point (which is very low). As a result, the Poisson model does not appear to be suitable for modelling the number of points per possession since it does not capture the actual scoring pattern in basketball.
Finally, when considering external validation Criteria 1–4, the Poisson model performs similarly to the normal RAPMs, being slightly worse in Criteria 1–3 but better in 4 out of 6 box scores used for Criterion 4. However, when compared to the multinomial model, the Poisson model performs worse in relation to the box scores for Criterion 4: the multinomial model outperforms the Poisson in 4 out of 6 criteria, while the Poisson is better in only one (blocks), and both models perform equally on offensive rebounds.
To conclude, the Poisson model is not compatible with the possible outcomes for the number of points per possession modeled in this work. Its final probability distribution fails to capture the non-decreasing pattern of probabilities observed in basketball, and the Poisson model performs worse than the multinomial model in external validation criteria based on box scores. Thus we cannot propose its use for estimating players contributions in per-possession play-by-play data.
Figure 7: Estimated probabilities for each scoring category obtained using the Poisson model. The narrow, light gray bars represent the empirical frequencies of points per possession (in the category of 4, 5 and 6 points the frequency is quite low).
4.6 Ordered Multinomial Model implementation
The ordered multinomial regression serves as an alternative to the multinomial model, treating the response variable as ordinal. However, the ordered multinomial regression relies on the important assumption of proportional odds. This assumption states that the relationship between each type of points scored per possession666measured by the differences in log-odds of cumulative scoring probabilities for subsequent types remains the same across different specifications of the team lineups. For instance, the log-odds of scoring two points or less, minus the log-odds of scoring one point or no points, should be constant regardless of the team lineups.
However, this assumption is clearly violated in per-possession basketball data, as each player’s efficiency varies across different types of points. Some players are exceptional at three-pointers, while others are more effective close to the basket, making them more accurate at scoring two-pointers. This inconsistency is evident in Figure 8, where the variation of the players’ coefficients across different scoring categories indicates a violation of the proportional odds assumption, suggesting that players’ effects are not consistent across the levels of the response variable. The violation of the proportional odds assumption is further confirmed by the Brant test (p-value $<$ 0.001) implemented for the non-regularized ordinal logistic model. Finally, the ordinal model appears to have very close performance with regard to the external validation Criteria 1–4.
Figure 8: Log-odds of cumulative probabilities across the linear predictor for each scoring category.
@@ 5.1 The metric
In this section, we proceed by introducing an improved weighted version of EPTS which takes into consideration the participation of each player throughout the season. The goal of this evaluation metric is to quantify the expected points per possession for the team of the player under study. We will evaluate two different lineup configurations: one lineup including the player of interest and a second one without the player of interest. All other players on the court are assumed to belong to the reference category. The corresponding expected points of these two line-ups will be weighted according to the proportion of possessions in which the player of interest participated.
@@
As referee suggested, the weighted EPTS resembles the method for calculating expected values in the mixture framework. However, this literature discusses the interpretation of expected values computed as a weighted mean of zeros and positive values. In fact, the expected values for the quantitative outcome consistently show a downward shrinkage. In our context, shrinkage to zero is desirable because we aim to estimate the overall contribution of players to their team. This contribution should be influenced by the amount of time a player spends in the game. Players with limited playing time (e.g., LTPs) are unlikely to have a significant impact on their team, even if their performance during play is exceptional. Therefore, EPTS and wEPTS can be used complementary: EPTS can identify a limited number of players with “positive” performance even with low playing time, while wEPTS measures overall contribution, which accounts for playing time. By using wEPTS—a weighted mean of zero (when the player is not playing) and their mean points (when playing)—we provide a robust and meaningful measure of player contribution that addresses the LTP issue.
@@ 5.2 Weighted EPTS (wEPTS) through validation criteria
The RAPM and EPTS ratings presented in the previous sections are based on biased “shrunk” estimates of model coefficients which quantify the contribution of each player. This section explores the impact of bias on RAPM and EPTS ratings. We compare two approaches: (a) using directly biased coefficients derived using lasso and (b) using the unbiased (MLE) coefficients, after removing players whose coefficients were flagged as zero by lasso. For the latter case we will use the conventional name “after-lasso” ratings.
@@
5.4 The Effect of Low Time Players
In Section 3.3.3, the implementation of a cut-off threshold for minutes played by the studied players is examined in relation to the higher performance of the regularized Gaussian models with respect to players ranking. Based on these improved results, we subsequently applied the logistic multinomial model to the “filtered” dataset, excluding low-time players from the analysis. Nevertheless, it would be worthwhile to examine the proposed methodology by considering all players in the analysis.
In fact, we expect coming across similar issues regarding the low-time players (LTPs) during the development of the EPTS; however, we do not expect such issues to arise with the wEPTS. Our intuition is confirmed, as shown in Figure 9, where the distribution of the top EPTS contributed players is right skewed when LTPs are incorporating in the analysis, and more flat when they are excluded; additionally, LTPs (less than 200 minutes in the season) constitute 38% of top 100 in the full data analysis, indicating that LTPs have a significant impact on the results. As for the wEPTS, the corresponding distribution is similar for both implementations. Hence, wEPTS appears to address the issue of LTPs, as the distribution of the top 100 remains robust despite the presence of LTPs.
Figure 9: Density of top-100 contributed players playing time based on the Multinomial approach followed by including (Full dataset) or excluding low-time players (Filtered dataset).
6 Concluding remarks
6.1 Conclusion
Following the literature, we initially implemented ridge regression on the full possession dataset for the 2021-2022 NBA season. The RAPM (Regularized Adjusted Plus-Minus) model is generally applied to data aggregated over shifts, where the composition of the team on the court remains constant. This aggregation lends itself to the use of Gaussian models. However, applying a Gaussian model directly to disaggregated data presents significant challenges in terms of justification. Implementing a linear model in this context can lead to a substantial oversimplification and may introduce bias, as has been observed. Specifically, this approach was susceptible to the influence of low-time players (LTPs), whose limited playing time inflated their estimated contributions, despite the shrinkage. To address this issue, we have investigated two different approaches: First, we have implemented lasso instead of ridge and, second, we have removed low-time players from RAPM estimation.
For the first approach, lasso regression promotes sparsity by setting specific coefficients equal to zero. This will be effective for less influential variables (players). This characteristic of lasso resulted in RAPM estimates with a clearer distinction between well-performed, average and low-performed players. Concerning LTPs, the problem was mitigated (but not diminished) by shrinking the coefficients of such players to zero (i.e. to the group of “average” players). On the other hand, the exclusion of LTPs (defined as less than 200 minutes for the season) also improved the performance of the RAPM ratings. For this reason, we have proceeded by combining the two strategies (i.e. implementing lasso without LTPs).